# Exploring Holomorphic Retracts

### v1

@@ 2.2. Notions of extremality and strict convexity
Proof of Claim-1: If possible assume that $a\in M$. Choose $s_{0},~{}t_{0}\in\mathbb{R}$ such that $|s_{0}|<\epsilon$ and $|s_{0}t_{0}|<\epsilon$. Since $a\in M$, we get $a+s_{0}u\in M$. Hence we get $a+s_{0}u=ca$ for some $c\in\mathbb{R}$. Taking norm on both sides we get $|c|=1$.
$\|a+s_{0}t_{0}u\|^{2}=\|(1-t_{0})a+t_{0}(a+s_{0}u)\|^{2}=(1-t_{0})^{2}+2t_{0}(1-t_{0})c+c^{2}t_{0}^{2}<(1-t_{0})^{2}+2t_{0}(1-t_{0})+t_{0}^{2}=(1-t_{0}+t_{0})^{2}=1$.
Hence we get $\|(1-t_{0})a+t_{0}(a+s_{0}u)\|^{2}<1$. This implies that $\|a+s_{0}t_{0}u\|<1$. Hence, we get $a+s_{0}t_{0}u\notin\partial\Omega$. This contradicts our assumption, proving the claim.
By Hahn–Banach theorem, there exists a $\mathbb{C}$- linear functional $f$ with $\|f\|=1$ such that $f(z)=0$ for all $z\in M$ and $f(a)=dist(a,M)$.
Claim-2: ${\rm dist}(a,M)=1$.
Proof of Claim-2: Let us begin by recalling that $\rm dist(a,M)=\inf\{\|a+su\|:s\in\mathbb{R}\}$ and note firstly, that ${\rm dist}(a,M)\leq 1$. Take an arbitrary $s\in\mathbb{R}$. Choose $t_{0}\in[0,1)$ such that $|(1-t_{0})s|<\epsilon$ and note that
$\|(1-t_{0})(a+su)+t_{0}a\|\leq(1-t_{0})\|a+su\|+t_{0}\|a\|=(1-t_{0})\|a+su\|+t_{0}.$
We simplify the above inequality to write it as
(2.1) $\|a+(1-t_{0})su\|\leq(1-t_{0})\|a+su\|+t_{0}.$
Since $|(1-t_{0})s|<\epsilon$, we get from (2.1) that $\|a+(1-t_{0})su\|=1$. This implies $\|a+su\|\geq 1$ for all $s\in\mathbb{R}$. Hence, $dist(a,M)=1$, finishing the proof of claim-2.
Note that $f(a+tu)=f(a)=1$. Since $\Omega$ is strictly $\mathbb{C}$-convex, we have $a=a+tu$. This implies that $u=0$. So $\Omega$ is $\mathbb{R}$-convex. ∎
@@
The purpose of recalling this notion in this article about retracts is the following result; being a result about retracts, we prefer to include the details of the proof.
Theorem 2.20.
(A slight modification of proposition $4.5$ in [JJ]). Let $D$ be any bounded balanced pseudoconvex domain in $\mathbb{C}^{N}$. Suppose $Z$ is a non-trivial retract through the origin with the property that the boundary points of $D_{L}:=D\cap T_{0}Z$ viewed as a domain in $T_{0}Z$ are holomorphically extreme. Then, $Z$ is linear.
Proof.
Let $\mathfrak{h}_{D_{L}}$ denote the Minkowski function on $D_{L}$. Since $Z$ is a retract of $D$, there exists a holomorphic retraction map $\rho:D\to D$ such that $\rho(D)=Z$. Note that $\rho|_{D_{L}}$ is a holomorphic map from $D_{L}$ to $D$ and $Z=image(\rho|_{D_{L}})$. Observe that $Z$ is a connected complex submanifold and complex tangent space at 00 is $T_{0}Z=\{v\in\mathbb{C}^{N}:\rho^{\prime}|_{D_{L}}(0)v=v\}$. Let $v\in D_{L},v\neq 0$. By our assumption on $\partial D_{L}$, it follows that the mapping
$\phi:\Delta\ni\lambda\longmapsto\lambda v/\mathfrak{h}_{D_{L}}(v)\in D_{L}.$
is the unique (modulo $Aut(\Delta)$) $K_{D_{L}}$ geodesic for $(0,v/\mathfrak{h}_{D_{L}}(v))$. Since $D_{L}\subset D$ and $L$ is a linear subspace of $\mathbb{C}^{N}$, $\mathfrak{h}_{D_{L}}(v)=\mathfrak{h}_{D_{L}}(v)$. So $\phi$ is the unique (modulo $Aut(\Delta)$) $K_{D}$ geodesic for $(0,v/\mathfrak{h}_{D_{L}}(v))$. Note that
$K_{D}(\rho|_{D_{L}}\circ\phi(0);(\rho|_{D_{L}}\circ\phi)^{\prime}(0))=K_{D}(0,v/\mathfrak{h}_{D_{L}}(v))=K_{D}(\phi(0);\phi^{\prime}(0)).$
Hence $\rho|_{D_{L}}\circ\phi$ is a geodesic for $(0,v/\mathfrak{h}_{D_{L}}(v))$. By uniqueness, it follows that $\rho|_{D_{L}}\circ\phi=\phi\circ\psi$, where $\psi\in Aut(\Delta)$. Substituting $\psi^{-1}(\mathfrak{h}_{D_{L}}(v))$ which is a point in $\Delta$, into the above equation, we see that $v$ lies in the image of $\rho|_{D_{L}}$. Hence $D_{L}\subset Z$. Now $T_{0}Z$ and $Z$ have the same dimension. Since $Z$ is connected, it follows from the principle of analytic continuation that $Z=D_{L}$. Hence $Z$ is linear. ∎
@@ 2.3. Geodesics and totally geodesic subsets
It follows atleast for bounded convex domains $D\subset\mathbb{C}^{N}$ that retracts of $D$ are totally geodesic, a notion that shall be made use of, while recalling the work [Mok] of Mok.
@@ 2.4. One dimensional Retracts
Similarly we can define complex $K$-geodesic (or complex Kobayashi geodesic) if we replace $c_{D}$ by $K_{D}$ in the above definition. By Lemma 2.6.1 (ii) i [Abate_Itrtn], every complex $K$-geodesic is a complex $C$-geodesic. But the converse is not true because there are domains in $\mathbb{C}^{N}$ which are Kobayashi hyperbolic but not Caratheodory hyperbolic. Owing to the direct relationship between $C$-geodesics and retracts, we shall mostly work with $C$-geodesics rather than $K$-geodesics.
E. Vesentini [Vesentini_1], gives the following characterization of the complex geodesics of $D$.
@@
From proposition 2.10, we can conclude that not all linear subspaces of bounded balanced domain $D$ are a retract of $D$ unless $D$ is convex. One natural question arises: which linear subspaces of bounded balanced domain $D$ does not arises as a retract of $D$. The following corollary gives more general result. Before going to the statement, first we recall the notion of convexity at a boundary point.
@@
Let $D=\{z\in\mathbb{C}^{N}:\mathfrak{h}(z)<1\}$ be a balanced domain in $\mathbb{C}^{N}$ ($\mathfrak{h}$ is the Minkowski function on $D$) and $a\in D$. $D$ is said to be convex at $\frac{a}{\mathfrak{h}(a)}$ if there exists a $\mathbb{C}$- linear functional $L:\mathbb{C}^{N}\to\mathbb{C}$ with $|L|\leq\mathfrak{h}$ and $|L(a)|=\mathfrak{h}(a)$.
@@
Let $D$ be a bounded balanced pseudoconvex domain in $\mathbb{C}^{N}$ and $a\in D$ such that D is not convex at $a/\mathfrak{h}(a)$. Then there does not exists one dimensional retract of $D$ passing through 00 and $a$.
@@
Assume the contrary, there exists a one dimensional retract $R$ of $D$ passing through 00 and $a$. By theorem 2.25, $R$ is image of some complex C-geodesic $\varphi:\Delta\to D$. But by lemma 11.1.2 in [Jrncki_invrnt_dst] (see also remark 11.1.3(b) therein), there is no complex $C$- geodesic in $D$. This leads to a contradiction. ∎
@@ 2.5. Fixed point sets and Retracts
[…]
@@
Our goal here is to characterize the holomorphic retracts of the Cartesian product $B_{1}\times\ldots\times B_{N}$ (which will be referred to as polyballs for short), where the $B_{j}$’s are bounded balanced strictly convex domains or equivalently, the unit ball with respect to some norm in $\mathbb{C}^{n_{j}}$ for some $n_{j}\in\mathbb{N}$, for each $j=1,\ldots,N$; as noted in the previous section, this is equivalent to their boundaries being $\mathbb{R}$-extremal. As it turns out that the essential arguments for the case of a product of a pair of such balls extend to the more general case, we shall for simplicity of exposition, focus attention for the most part on the case of a product of $2$ such balls. So, let $B_{1}\subset\mathbb{C}^{N_{1}},B_{2}\subset\mathbb{C}^{N_{2}}$ (where $N_{1},N_{2}\in\mathbb{N}$ with say, $\min\{N_{1},N_{2}\}=N_{1}$) denote a pair of bounded balanced convex domains with $\mathbb{R}$-extremal boundaries; in-particular, $B_{1},B_{2}$ are balls with respect to some norms (on $\mathbb{C}^{N_{1}},\mathbb{C}^{N_{2}}$ respectively) and needless to say are domains of holomorphy. The simplest example of such a domain is the standard unit disc in $\mathbb{C}$ which we denote by $\Delta$. However, many nice properties of the disc may well not hold for such domains (i.e., bounded balanced convex domains). To begin with, balls in the $l^{p}$-norm for $1<p<\infty$, while satisfying the condition that their boundary points are $\mathbb{R}$-extremal, need not always be infinitely smooth – they are atmost $[p]$-smooth, where []delimited-[][$\cdot$][ ⋅ ] denotes the greatest integer function. They are thus atleast $C^{1}$-smooth and to get an example where this $C^{1}$-smoothness fails, we may take an intersection of a pair of suitable dilates of the ball in the $l^{p_{1}}$ norm with that of the $l^{p_{2}}$ norm, for a pair of positive distinct reals $p_{1},p_{2}$ in the interval $(1,\infty)$ to get examples of domains whose boundaries are not $C^{1}$-smooth but nevertheless satisfy the $\mathbb{R}$-extremal condition; as this is essentially the well-known fact that the intersection of a pair of strictly convex domains is again strictly convex, we shall provide details with $\mathbb{R}$-extremality replaced by $\mathbb{C}$-extermality, in the following.
@@
If $D_{1},D_{2}$ are domains in $\mathbb{C}^{N}$ with $\mathbb{C}$-extremal boundary, then $D:=D_{1}\cap D_{2}$ is also a domain with $\mathbb{C}$-extremal boundary.
@@
Let $a$ be an arbitrary point in $\partial D$. We want to prove that $a$ is $\mathbb{C}$-extremal point in $D$. If possible assume that $a$ is not $\mathbb{C}$-extremal in $D$. That is, there exists non-zero $u\in\mathbb{C}^{N}$ and there exists $\epsilon>0$ such that $a+tu\in\partial D~{}\text{for~{} all }~{}t\in\mathbb{C}~{}\text{with}~{}|t|<\epsilon$. In particular, $a\in\partial D$ and we shall split the proof into cases depending upon which ‘piece’ of the boundary $a$ lies; herein, piece refers to one of the components in the decomposition of the boundary of $D$ as
@@
Without loss of generality assume that $\epsilon=1$. Consider a map $\alpha:\Delta\to\mathbb{C}$ defined by $\alpha(t)=a+tu$. Since $u\neq 0$, $\alpha$ is a non-constant holomorphic map. Note that $\alpha$ is a homeomorphism from $\Delta$ onto $\alpha(\Delta)$. Consider the sets $S_{1}:=\{t\in\Delta:\alpha(t)\in\partial{D_{1}}\}$ and $S_{2}:=\{t\in\Delta:\alpha(t)\in\partial{D_{2}}\}$. By above decomposition, $S_{1}\cup S_{2}=\Delta$ and note that $S_{1}=\alpha^{-1}(\partial D_{1})$ and $S_{2}=\alpha^{-1}(\partial D_{2})$. Again by continuity of $\alpha$, $S_{1}$ and $S_{2}$ are closed sets in $\Delta$.
@@
So we get $S_{2}=\Delta$. Hence we can conclude that $S_{1}$ or $S_{2}$ has atleast one interior point.
Without loss of generality assume that $S_{1}$ has an interior point. Let $t_{0}\in\Delta$ be an interior point of $S_{1}$. This means that there exists $\delta>0$ such that $A:=\{t\in\mathbb{C}:|t-t_{0}|<\delta\}\subset S_{1}$. Choose any $s\in\mathbb{C}$ with $|s|<\delta$. Let $b:=a+t_{0}u$. Note that $b\in\partial D_{1}$, because $t_{0}\in S_{1}$. We can write $b+su=a+(t_{0}+s)u$ and observe that $t_{0}+s\in A$. Hence $t_{0}+s\in S_{1}$ for all $s\in\mathbb{C}$ with $|s|<\delta$. By definition of $S_{1}$, we obtain that $b+su\in\partial D_{1}$ for all $s\in\mathbb{C}$ with $|s|<\delta$. Since $D_{1}$ is $\mathbb{C}$-extremal, we have $u=0$. This contradicts our assumption that $a$ is not a $\mathbb{C}$-extremal point of $D$. Hence we conclude that $a$ is $\mathbb{C}$-extremal. ∎

### v6

@@ 2.2. Notions of extremality and strict convexity
Proof of Claim-1: If possible assume that $a\in M$. Since $a\in\partial\Omega$, there exists non-zero $\lambda\in\mathbb{C}$ such that $a=\lambda u$. Choose $t_{0}$ be a non-zero real number such that $|t_{0}|<\epsilon$ and $|\lambda+t_{0}|\neq|\lambda|$. Taking norm on both sides of the equation $a=\lambda u$, we get $|\lambda|=1/\|u\|$. From the assumption $a+tu\in\partial\Omega$, we obtain
$1=\|a+t_{0}u\|=\|\lambda u+t_{0}u\|=\|u\|\cdot|\lambda+t_{0}|=\frac{|\lambda+t_{0}|}{|\lambda|}.$
The above equations tells that $|\lambda+t_{0}|=|\lambda|$. This contradicts our choice of $t_{0}$ and finishes the proof of the claim.
[…]
Proof of Claim-2: Let us begin by recalling that ${\rm dist}(a,M)=\inf\{\|a-su\|:s\in\mathbb{C}\}$ and note firstly (by considering $s=0$) that ${\rm dist}(a,M)\leq 1$. To prove this claim then, it suffices to show that $\rm dist(a,M)\geq 1$. To this end, it is convenient to rewrite the definition of the aforementioned distance as ${\rm dist}(a,M)=\inf\{\|a+su\|:s\in\mathbb{C}\}$ – this is not a mere sign change, it is a changing over to the ‘dual’ viewpoint that the aforementioned distance is also the infimum of distances of the origin to points on the affine line through $a$ in the direction of $u$. Since $D$ is convex, there exists a supporting hyperplane for every $p\in\partial D$. In particular, there exists a supporting hyperplane $\widetilde{H}=\{z\in\mathbb{C}^{N}:g(z)=g(a)\}$ (for some $\mathbb{R}$-linear functional $g$) at the point $a$. Let $l:=\{a+tu:t\in\mathbb{R}\}$. Let us first verify that $l\subset\tilde{H}$. To prove this by contradiction, assume the contrary that $l\not\subset\tilde{H}$. Consider the linear subspace $H:=\{z\in\mathbb{C}^{N}\;:\;g(z)=0\}$. By the definition of supporting hyperplane, $a\notin H$. So, $span\{a\}\cap H=\{0\}$.
We now show that $span\{u\}\cap H=\{0\}$ as well. Indeed, if $\lambda u\in H$ for some $\lambda\in\mathbb{C}$, then $g(\lambda u)=\lambda g(u)=0$. Since $l\not\subset\widetilde{H}$, there exists a non-zero real number $t_{0}$ such that $a+t_{0}u\notin\widetilde{H}$. This means that $u\notin\ker(g)$. From $\lambda g(u)=0$, we obtain that $\lambda=0$. Hence $span\{u\}\cap H=\{0\}$.
Now by Claim-1, $a$ and $u$ are linearly independent. Since $span\{a\}\cap H=\{0\}$ and $span\{u\}\cap H=\{0\}$, it follows that $dim(H)<N-1$. This contradicts the fact that $H$ is a hyperplane. Hence $a+su\in\widetilde{H}$ for all $s\in\mathbb{R}$. Therefore $\|a+su\|\geq 1$. This implies that $\rm dist(a,M)\geq 1$. This completes the proof of the claim.
Completing the proof of the theorem consists in now observing that: by the Hahn–Banach theorem, there exists a $\mathbb{C}$-linear functional $f$ with $\|f\|=1$ such that $f(z)=0$ for all $z\in M$ and $f(a)=\rm dist(a,M)$. Note then that $f(a+tu)=f(a)=1$. Since $\Omega$ is strictly $\mathbb{C}$-convex, this means $a=a+tu$, which in turn implies $u=0$. So $\Omega$ is $\mathbb{R}$-convex, finishing the proof in all. ∎
Remark 2.22.
We shall henceforth not explicitly mention the field and simply refer to a strictly $K$-convex domain as strictly convex domain.
@@
The purpose of recalling this notion in this article about retracts is to record the details for proposition 1.4 in the introductory section; which will be done in the next section.
@@ 2.3. Geodesics and totally geodesic subsets
It follows atleast for bounded convex domains $D\subset\mathbb{C}^{N}$ that retracts of $D$ are totally geodesic, a notion that shall be made use of, while recalling the work [Mok] of Mok. This lemma also shows that if $X$ is hyperbolic then every retract of $X$ is hyperbolic.
@@ 2.4. One dimensional Retracts
Here $c_{D}$ denotes the Caratheodory distance; as we shall be working with a variant of this as well, we recall both of them and their notations:
$c_{D}(z,w)=\sup\{p(f(z),f(w)):f\in\mathcal{O}(G,\Delta)\}$
The variant of $c_{D}$ denoted $c_{D}^{*}$ which will also be useful later for us, is defined as follows.
$c_{D}^{*}(z,w)=\sup\{m(f(z),f(w)):f\in\mathcal{O}(G,\Delta)\}$
‘Dualizing’ this definition i.e., by swapping the roles of $G$ and $\Delta$, considering $\mathcal{O}(\Delta,G)$ in place of $\mathcal{O}(G,\Delta)$, we have what is known as the Lempert function of $D$ denoted $l_{D}$. For the precise definition of this invariant function as well as other invariant metrics and distances alongwith various of their fundamental properties particularly of the Kobayashi metric, we refer to the best known book on the subject [Jrncki_invrnt_dst]; here we shall only recall to the extent of relevance to retracts and convenience of later use in this article. Thus, we only mention that we can similarly define complex $K$-geodesic (i.e., complex Kobayashi geodesic) if we replace $c_{D}$ by $K_{D}$ in the above definition. By Lemma 2.6.1 (ii) in [Abate_Itrtn], every complex $C$-geodesic is a complex $K$-geodesic. But the converse is not true because there are domains in $\mathbb{C}^{N}$ which are Kobayashi hyperbolic but not Caratheodory hyperbolic. Owing to the direct relationship between $C$-geodesics and retracts, we shall mostly work with $C$-geodesics rather than $K$-geodesics. Vesentini [Vesentini_1], gives the following characterization of the complex geodesics of $D$.
@@
Recall that by proposition 2.11, not all one-dimensional linear subspaces of bounded balanced domain $D$ are retracts of $D$, unless $D$ is convex. This naturally leads to the question about which are those one-dimensional linear subspaces of bounded balanced non-convex domain $D$ which fail to be a retract of $D$? We next lay down a corollary that answers this; its statement requires the notion of convexity at a boundary point, whose definition we recall from [Jrncki_invrnt_dst].
@@
Let $D=\{z\in\mathbb{C}^{N}:h(z)<1\}$ be a balanced domain in $\mathbb{C}^{N}$ where as usual $h$ denotes the Minkowski functional of $D$. Let $a$ be any point of $D$, other than the origin such that $D$ is bounded in the direction of $a$; note that this forces $h(a)$ to be non-zero. The domain $D$ is said to be convex at the boundary point $a/h(a)$, if there exists a $\mathbb{C}$-linear functional $L:\mathbb{C}^{N}\to\mathbb{C}$ with $|L|\leq h$ and $|L(a)|=h(a)$. Note that the latter (equality) condition may be written as $|L(p)|=1$ where $p=a/h(a)\in\partial D$. When the aforementioned pair of conditions hold, we also say that $\partial D$ is convex at the boundary point $p$ (or that the balanced domain $D$ is convex near the boundary point closest in the direction of $a$). The boundary $\partial D$ (or the domain $D$ itself) is said to be convex near $a/h(a)$ if there exists an open ball $B$ around $a/h(a)$ such that $B\cap\partial D$ is convex at every $q\in B\cap\partial D$.
One may want to relate the above notion of convexity with the more standard geometric one whose requirement does not involve its boundary points explicitly. Namely, we recall that $D$ is said to be geometrically convex near the boundary point $p$ if there exists an open set $U$ about $p$ such that $U\cap D$ is convex in the usual sense that for every pair of points $p_{1},p_{2}\in U\cap D$, we have $(1-t)p_{1}+tp_{2}\in U\cap D$ for all $t\in[0,1]$.
The relation between the above pair of notions of convexity will be discussed in section 6. Here we shall clarify their connection with another standard one namely, the analogue of the above definition 2.30 obtained by restricting to real scalars instead of complex ones.
Definition 2.31.
Let $D=\{z\in\mathbb{C}^{N}:h(z)<1\}$ be a balanced domain in $\mathbb{C}^{N}$ where as before $h$ denotes the Minkowski functional of $D$. Let $a$ be any point of $D$, other than the origin such that $D$ is bounded in the direction of $a$ i.e., $h(a)\neq 0$. The domain $D$ is said to be convex at the boundary point $a/h(a)$, if there exists a $\mathbb{R}$-linear functional $f:\mathbb{C}^{N}\to\mathbb{R}$ with $|f|\leq h$ and $|f(a)|=h(a)$. We note that this definition can be made for any real vector space, of-course $\mathbb{R}^{n}$ being the case of our interest.
Proposition 2.32.
Let $D$ be a balanced domain in $\mathbb{C}^{N}$. The above two definitions 2.30 and 2.31 are equivalent.
Proof.
Assume that $D$ is convex at $a/h(a)$ in the sense of definition 2.30. Then there exists a $\mathbb{C}$-linear functional $L:\mathbb{C}^{N}\to\mathbb{C}$ such that $|L(z)|\leq h(z)$ for all $z\in\mathbb{C}^{N}$ and $|L(a)|=h(a)$. Consider the real linear functional $:\mathbb{C}^{N}\to\mathbb{R}$ defined by $T(z)=Re(\alpha L(z))$, where $\alpha=|L(a)|/L(a)$. Note that $|\alpha|=1$ and we obtain that
$|T(z)|=|Re(\alpha L(z))|\leq|\alpha||L(z)|=|L(z)|\leq h(z).$
Hence we obtain that $|T(z)|\leq h(z)$ for all $z\in\mathbb{C}^{N}$ and $|T(a)|=Re(\alpha L(a))=|L(a)|=h(a)$. Therefore $D$ is convex at $a/h(a)$ according to definition 2.31.
Conversely, assume that $D$ is convex at $a/h(a)$ as per definition 2.31. Then there exists a $\mathbb{R}$-linear functional $f:\mathbb{C}^{N}\to\mathbb{R}$ with $|f(z)|\leq h(z)$ and $|f(a)|=h(a)$. Define a $\mathbb{C}$-linear functional $L:\mathbb{C}^{N}\to\mathbb{C}$ by $L(z)=f(z)-if(iz)$. Observe that for each $z\in\mathbb{C}^{N}$, there exist $\lambda_{z}\in\mathbb{C}$ with $|\lambda_{z}|=1$ such that $\lambda_{z}L(z)=|L(z)|$. Note that
$|L(z)|=|Re(\lambda_{z}L(z))|=|Re(L(\lambda_{z}z))|=|f(\lambda_{z}z)|\leq h(\lambda_{z}z)=|\lambda_{z}|h(z)=h(z)$
Hence we get that $|L(z)|\leq h(z)$ for all $z\in\mathbb{C}^{N}$. In-particular, $|L(a)|\leq h(a)$. But then the very definition of $L$ shows that $|L(a)|\geq h(a)$ and we thereby get that $|L(a)|=h(a)$ (and $ia$ lies in the kernel of $f$ though this observation is not necessary for the proof per se). Hence we have proved that $D$ is convex at $a/h(a)$ according to definition 2.30. ∎
The next pair of results show that the existence (or non-existence) of retracts of a balanced pseudoconvex domain through the origin in any given direction is in essence, completely determined by the convexity of the nearest boundary point in that direction, as clarified by the foregoing definitions.
@@
Let $D$ be a balanced pseudoconvex domain in $\mathbb{C}^{N}$ and $a\in D$ such that $D$ is not convex at $a/h(a)$. Then there does not exist any one-dimensional retract of $D$ passing through 00 and $a$.
@@
Assume the contrary, there exists a one-dimensional retract $R$ of $D$ passing through 00 and $a$. By theorem 2.29, $R$ is image of some complex C-geodesic $\varphi:\Delta\to D$. But by lemma 11.1.2 in [Jrncki_invrnt_dst] (see also remark 11.1.3(b) therein), there is no complex $C$-geodesic in $D$. This leads to a contradiction. ∎
Proposition 2.34.
Let $D$ be a balanced pseudoconvex domain in $\mathbb{C}^{N}$ and $a\in D$ such that $D$ is convex at $a/h(a)$. Then $\varphi:\Delta\to D$ defined by $\varphi(\lambda)=\lambda a/h(a)$ is a complex $C$-geodesic on $D$. In particular, there always exists a one dimensional linear retract of $D$ passing through 00 and $a$.
Proof.
Let $\lambda$ be any non-zero element in $\Delta$. Since $D$ is convex at $a/h(a)$ and by proposition 2.3.1(b) in [Jrncki_invrnt_dst], we get
$c_{D}^{*}(\varphi(0),\varphi(\lambda))=c_{D}^{*}(0,\lambda a/h(a))=h(\lambda a/h(a))=|\lambda|$
Note that $c_{D}=\tanh^{-1}c_{D}^{*}$. So
$c_{D}(\varphi(0),\varphi(\lambda))=\tanh^{-1}\left(c_{D}^{*}(\varphi(0),\varphi(\lambda))\right)=\tanh^{-1}(|\lambda|)=p(0,\lambda)$
By proposition 11.1.4 in [Jrncki_invrnt_dst], $\varphi$ is a complex $C$-geodesic. ∎
We spell out the following immediate consequence as it settles a first basic question about the very existence of any non-trivial retract in bounded balanced domains of holomorphy of central interest of this article.
Proposition 2.35.
If $D$ is an arbitrary bounded balanced pseudoconvex domain in $\mathbb{C}^{N}$, then there exists atleast one (non-trivial) retract passing through origin.
Proof.
Let $a\in\partial D$ be the farthest point from the origin. Note that $D$ is convex at $a$. By the above proposition, there always exists a one-dimensional linear retract of $D$ passing through $a$. Alternatively, we may observe that the orthogonal projection onto ${\rm span}(a)$ maps $D$ onto it and thereby a retraction map on $D$. The aforementioned retract is the image of this retraction map. ∎
@@ 2.5. Fixed point sets and Retracts
Let $D$ be a bounded taut domain in $\mathbb{C}^{N}$. Assume that $f:D\to D$ be a holomorphic map which is not an automorphism and $\text{Fix}(f)$ is a non-empty subset of $D$ which consists of more than one point. Then there exists a non-trivial retract $Z$ of $D$ such that ${\rm Fix}(f)\subset Z$
@@
Our goal here is to characterize the holomorphic retracts of the Cartesian product $B_{1}\times\ldots\times B_{N}$ (which will be referred to as polyballs for short), where the $B_{j}$’s are bounded balanced strictly convex domains or equivalently, the unit ball with respect to some norm in $\mathbb{C}^{n_{j}}$ for some $n_{j}\in\mathbb{N}$, for each $j=1,\ldots,N$; as noted in the previous section, this is equivalent to their boundaries being $\mathbb{R}$-extremal. As it turns out that the essential arguments for the case of a product of a pair of such balls extend to the more general case, we shall for simplicity of exposition, focus attention for the most part on the case of a product of $2$ such balls. So, let $B_{1}\subset\mathbb{C}^{N_{1}},B_{2}\subset\mathbb{C}^{N_{2}}$ (where $N_{1},N_{2}\in\mathbb{N}$ with say, $\min\{N_{1},N_{2}\}=N_{1}$) denote a pair of bounded balanced convex domains with $\mathbb{R}$-extremal boundaries; in-particular, $B_{1},B_{2}$ are balls with respect to some norms (on $\mathbb{C}^{N_{1}},\mathbb{C}^{N_{2}}$ respectively) and needless to say are domains of holomorphy. The simplest example of such a domain is the standard unit disc in $\mathbb{C}$ which we denote by $\Delta$. However, many nice properties of the disc may well not hold for such domains (i.e., bounded balanced convex domains). To begin with, balls in the $l^{p}$-norm for $1<p<\infty$, while satisfying the condition that their boundary points are $\mathbb{R}$-extremal, need not always be infinitely smooth – they are atmost $[p]$-smooth, where []delimited-[][$\cdot$][ ⋅ ] denotes the greatest integer function. They are thus atleast $C^{1}$-smooth and to get an example where this $C^{1}$-smoothness fails, we may take an intersection of a pair of suitable dilates of the ball in the $l^{p_{1}}$ norm with that of the $l^{p_{2}}$ norm, for a pair of positive distinct reals $p_{1},p_{2}$ in the interval $(1,\infty)$ to get examples of domains whose boundaries are not $C^{1}$-smooth but nevertheless satisfy the $\mathbb{R}$-extremal condition; as this is essentially the well-known fact that the intersection of a pair of strictly convex domains is again strictly convex, we shall provide details with $\mathbb{R}$-extremality replaced by $\mathbb{C}$-extremality, in the following.
@@
If $D_{1},D_{2}$ are any pair of domains in $\mathbb{C}^{N}$ with weakly (resp. strongly) $\mathbb{C}$-extremal boundary, then so is every connected component of $D:=D_{1}\cap D_{2}$.
@@
Let $D_{1},D_{2}$ be pair of domains with weakly $\mathbb{C}$-extremal boundary. To prove this result by contradiction, assume $a\in\partial D$ is not weakly $\mathbb{C}$-extremal. That is, there exists non-zero $u\in\mathbb{C}^{N}$ and there exists $\epsilon>0$ such that $a+tu\in\partial D~{}\text{for~{} all }~{}t\in\mathbb{C}~{}\text{with}~{}|t|<\epsilon$. We shall split the proof into cases depending upon which ‘piece’ of the boundary $a$ lies; herein, piece refers to one of the components in the decomposition of the boundary of $D$ as
@@
Without loss of generality assume that $\epsilon=1$. Consider a map $\alpha:\Delta\to\mathbb{C}$ defined by $\alpha(t)=a+tu$. Since $u\neq 0$, $\alpha$ is a non-constant holomorphic map. Note that $\alpha$ is a homeomorphism from $\Delta$ onto $\alpha(\Delta)$, which is in $\partial D$. Consider the sets $S_{1}:=\{t\in\Delta:\alpha(t)\in\partial{D_{1}}\}$ and $S_{2}:=\{t\in\Delta:\alpha(t)\in\partial{D_{2}}\}$. By above decomposition, $S_{1}\cup S_{2}=\Delta$ and note that $S_{1}=\alpha^{-1}(\partial D_{1})$ and $S_{2}=\alpha^{-1}(\partial D_{2})$. Note that $\partial D_{i}\cap\partial D$ is a closed subset of $\partial D$ for $i=1,2$. By continuity of $\alpha$, $S_{1}$ and $S_{2}$ are closed sets in $\Delta$.
@@
So, we get $S_{2}=\Delta$. Hence we can conclude that $S_{1}$ or $S_{2}$ has atleast one interior point.
Without loss of generality assume that $S_{1}$ has an interior point. Let $t_{0}\in\Delta$ be an interior point of $S_{1}$. This means that there exists $\delta>0$ such that $A:=\{t\in\mathbb{C}:|t-t_{0}|<\delta\}\subset S_{1}$. Choose any $s\in\mathbb{C}$ with $|s|<\delta$. Let $b:=a+t_{0}u$. Note that $b\in\partial D_{1}$, because $t_{0}\in S_{1}$. We can write $b+su=a+(t_{0}+s)u$ and observe that $t_{0}+s\in A$. Hence $t_{0}+s\in S_{1}$ for all $s\in\mathbb{C}$ with $|s|<\delta$. By definition of $S_{1}$, we obtain that $b+su\in\partial D_{1}$ for all $s\in\mathbb{C}$ with $|s|<\delta$. Since $D_{1}$ is weakly $\mathbb{C}$-extremal, we have $u=0$. This contradicts our assumption that $a$ is not a weakly $\mathbb{C}$-extremal point of $\partial D$. Hence we conclude that $a$ is weakly $\mathbb{C}$-extremal.
Next, assume that $D_{1},D_{2}$ are domains in $\mathbb{C}^{N}$ with strongly $\mathbb{C}$-extremal boundary. Let $a\in\partial(D_{1}\cap D_{2})$. If possible assume that there exists $\epsilon>0$ and $u\in\mathbb{C}^{N}$ such that $a+tu\in\partial(D_{1}\cap D_{2})$ for all $t\in\mathbb{C}$ with $|t|<\epsilon$. Note that $\partial D_{1}\cap\partial D_{2}\subset\partial(D_{1}\cap D_{2})$. Hence $\overline{D_{1}\cap D_{2}}=\overline{D_{1}}\cap\overline{D_{2}}$. If $a+tu\in\overline{D_{1}\cap D_{2}}$ for all $t\in\mathbb{C}$ with $|t|<\epsilon$, then $a+tu\in\overline{D_{1}}\cap\overline{D_{2}}$ for all $t\in\mathbb{C}$ with $|t|<\epsilon$. Since $D_{1}$ and $D_{2}$ are domains with strongly $\mathbb{C}$-extremal boundary, we have $u=0$. Therefore, $D_{1}\cap D_{2}$ is a domain with strongly $\mathbb{C}$-extremal boundary. ∎